OpenStudy (anonymous):

Solve (x + 4) – (7x – 2) = 3.

OpenStudy (anonymous):

x+4 - 7x + 2 = 3 -6x + 6 = 3 Can you solve it now?

OpenStudy (anonymous):

it still aint givin me the right answer

OpenStudy (anonymous):

Take 6 common. 6( -x + 1 ) = 3

OpenStudy (anonymous):

(x + 4) – (7x – 2) = 3 x + 4 - 7x + 2 = 3 x - 7x + 4 + 2 = 3 -6x + 6 = 3 -6x = 3 - 6 - 6x = -3 -6x/-6 = -3/-6 x = 1/2...

OpenStudy (anonymous):

ohh i already got the answer but thank you and im need some more help on these problems

OpenStudy (anonymous):

which?